# KAG: Boosting LLMs in Professional Domains via Knowledge Augmented Generation

### v1

@@ 2.2.3 Domain Knowledge Injection And Constraints
When openIE is applied to professional domains, irrelevant noise will be introduced. Previous research has shown that noisy and irrelevant corpora can significantly undermine the performance of LLMs[3, 5, 13]. It is a challenge to align the granularity of extracted information and domain knowledge. The domain knowledge alignment capabilities in KAG include: 1) Domain term and concept injection. We use an iterative extraction approach, First, we store domain concepts and terms with description in KG storage. Second, we extract all instances in the document through openIE, then we perform vector retrieval to obtain all possible concept and term sets $E_{d}$. Finally, we add $E_{d}$ to the extraction prompt and perform another extraction to obtain a set $E_{d}^{a}$ that is mostly aligned with the domain knowledge. 2) Schema-constraint Extraction. In the vertical professional domains, the data structure between multiple documents in each data source such as drug instructions, physical examination reports, government affairs, online order data, structured data tables, etc. has strong consistency, and is more suitable for information extraction with schema-constraint, structured Extraction also makes it easier to do knowledge management and quality improvement. For detailed information about knowledge construction based on Schema-constraint, please refer to the SPG [^1] Official site of SPG: https://spg.openkg.cn/en-US and OneKE[14]. This article will not introduce it in detail. It is worth noting that, as shown in figure 2, for the same entity type, such as Person, we can pre-define properties and relations such as name, gender, placeOfBirth, (Person, hasFather, Person), (Person, hasFriend, Person), and can also extract tripples directly such as (Jay Chou, spgClass, Person), (Jay Chou, constellation, Capricorn), (Jay Chou, record company, Universal Music Group) through openIE. 3) Pre-defined Knowledge Structures By Document Type. Professional documents such as drug instructions, government affairs documents, and legal definitions generally have a relatively standardized document structure. Each type of document can be defined as an entity type, and different paragraphs are different properties of the entity. Taking government affairs as an example, we can pre-define the GovernmentAffair EntityType and properites such as administrative divisions, service procedures, required materials, service locations, and target groups. The divided chunks are the values of different properties. If the user asks "What materials are needed to apply for housing provident fund in Xihu District?", you can directly take out the chunk corresponding to property required materials to answer the question, avoiding the possible hallucinations caused by LLM re-generation.
@@ 2.2.4 Mutual indexing between text chunk vectors and knowledge structures
KAG’s mutual indexing is a knowledge management and storage mechanism that conforms to the LLMFriSPG semantic representation. As is described in section 2.1, it includes four core data structures: 1) Shared Schemas are coarse-grained-types pre-defined as SPG Classes at project level, it includes EntityTypes, ConceptTypes, and EventTypes, they serve as a high-level categorization such as Person, Organization, GEOLocation, Date, Creature, Work, Event. 2) Instance Graph include all event and entity instances in $KG_{cs}$ and $KG_{fr}$. that is, the instances constructed through schema-free openIE or through schema-constraint strict structure construction are both stored in the instance graph. 3) Text Chunks are special entity node that conforms to the definition of the Chunk EntityType. 4) Concept Graph is the core component for knowledge semantic alignment. it consists of a series of concepts and concept relations. Concept nodes are also fine-grained-types of instances. Through relation prediction, instance nodes can be linked to concept nodes to obtain their fine-grained semantic types. , and two storage structures: 1) Graph Store. Store graph data structures in LPG databases, such as TuGraph, Neo4J. 2) Text Vector. Store text and vectors in a vector storage engine, such as ES, or the vector storage embedded in the LPG engine.
@@ 2.3 Logical Form Solver
In the process of solving complex problems, two key steps are involved: reasoning and retrieval. Disassembling question is a planning process to determine the next problem to be tackled. Reasoning includes retrieving information based on the disassembled question, inferring the answer to the question according to the retrieved results, or re-disassembling the sub-question when the retrieved content cannot answer the question. Retrieval is to find the content that can be used as reference for the original question or the disassembled subquestion.
Since the interaction between different modules is based on the vector representation of natural language, inaccuracies often occur. Inspired by logical form, which is often used in KGQA, we design an executable language with both reasoning and retrieval capabilities, which breaks the question down into multiple logical expressions, each of which can contain functions for retrieval or logical operations. The mutual-structure index described in section 2.2 makes it possible to implement the above process. KAG Solver is inherently guided by logical form, which adopts the collaborative question decomposition and reasoning mechanism of ReSP[15] and logical form Executor. As is show in Figure 6, the global question decomposition is based on the ReSP method that uses global evidence memory and local pathway memory to solve the context redundant and over-planning caused by multiple rounds of retrieval through daul-summarizer. Under the guidance of the knowledge structure, logical form decomposes the question into multiple subquestions and represents them with logical function symbols, then proceeds to reason and solve each subquestion accordingly, as shown in Algorithm 17. Section 2.3.1, 2.3.2 and 2.3.3 introduce logical form function for Planning, logical form for Reasoning and logical form for Retrieval respectively. In general, the proposed logical form language has the following three advantages:
  • The use of symbolic language enhanced the rigor of problem decomposition and solving.
  • Deep integration of KG structure and text chunk in the retrieval process
@@
2.3.1 Logical Form Function
@@
$Retrieval(s=s_{i}:type[name],p=p_{i}:edge,o=o_{i}:type[name],$
@@
$Sort(A,direction=min|max,limit=n)$
@@
$math_{i}=Math(expr)$,
@@
$Deduce(left=A,right=B,op=entailment|greater|less|equal)$
@@
$Output(A,B,...)$
@@
Retrieval. According to the the knowledge or information retrieved from SPO, s, p, o should not repeatedly appear multiple times in the same expression. Constraints can be applied to the s, p, o for querying. For multi-hop queries, multiple retrievals are required. When the current variable refers to a previously mentioned variable, the variable name must be consistent with the referenced variable name, and only the variable name needs to be provided. The entity type and name are only specified during the first reference.
Sort. Sort the retrieval results. $A$ is the variable name for the retrieved SPO ($s_{i}$, $o_{i}$, or $s.prop$, $p.prop$, $o.prop$). $direction$ specifies the sorting direction, where $direction=min$ means sorting in ascending order and "direction=max" means sorting in descending order. $limit=n$ indicates outputting the topN results.
Math. Perform calculations. $expr$ is in LaTeX syntax and can be used to perform calculations on the retrieved results (sets) or constants. $math_{i}$ represents the result of the calculation and can be used as a variable name for reference in subsequent actions.
@@ 2.3.2 Logical Form for Reasoning
When the query statement represented by natural language is applied to the search, the logic is often fuzzy, such as "find a picture containing vegetables or fruits" and "find a picture containing vegetables and fruits". Whether text search or vector search is used, the similarity between the two queries is very high, but the corresponding answers are quite different. The same is true for problems involving logical reasoning processes such as and or not, and intersection differences. To this end, we use logical form to express the question, so that it can express explicit semantic relations. Similar to IRCOT, we decompose complex original problems and plan out various execution actions such as multi-step retrieval, numerical reasoning, logical reasoning, and semantic deduce. Each sub-problem is expressed using logical form functions, and dependencies between sub-questions are established through variable references. The inference resolution process for each sub-question is illustrated as Algorithm 11. In this process, the ‘GraphRetrieval’ module performs structured retrieval based on the logical function, with the retrieval results also being structured as SPO (Subject-Predicate-Object) information. The ‘HybridRetrieval’ module, on the other hand, retrieves chunks based on the sub-question and the structured information. To understand how logical functions can be utilized to reason about complex problems, refer to the following examples as Table 2.
@@
Numerical Reasoning
Action1: Retrieval(s=s1:Player[Cristiano Ronaldo], p=p1:PlayedFor, o=o1:SportsTeam, p.PlayedForInYear=2011)
Step2: Determine the Foundation Years of Each Identified Team .
Action2: Retrieval(s=o1, p=p2:FoundationYear, o=o2:Year)
Step3: Which team was founded last ?
Action3: Sort(set=o1, orderby=o2, direction=max, limit=1)
Step1: What is the sum of 30 + 6 ?
Action1: math1 = Math(30+6)
Step2: Who is the founder of Tesla?
Action2: Retrieval(s=s2:Company[Tesla], p=p2:Founder, o=o2)
Step3: In which year was the founder of Tesla born?
Action3: Retrieval(s=o2, p=p3:YearOfBirth, o=o3)
Step4: How old will the founder of Tesla be in the year 2027?
Action4: math4 = Math(2027-o3)
Step5: What is the sum of math1 and math4?
Action5: math5 = Math(math1+math4)
Logical Reasoning
question: Find a picture containing vegetables or fruits.
Step1: Find pictures containing vegetables.
Action1: Retrieval(s=s1:Image, p=p2:Contain, o=o1:Vegetables)
Step2: Find pictures containing fruits.
Action2: Retrieval(s=s2:Image, p=p2:Contain, o=o2:Fruits)
Step3: Output s1, s2.
Action3: Output(s1, s2)
question: Find a picture containing vegetables and fruits.
Step1: Find pictures containing vegetables.
Action1: Retrieval(s=s1:Image, p=p2:Contain, o=o1:Vegetables)
Step2: Find pictures containing fruits.
Action2: Retrieval(s=s1, p=p2:Contain, o=o2:Fruits)
Step3: Output s1.
Action3: Output(s1)
Semantic Deduce
Step1: What documents are required to apply for a passport?
Action1: Retrieval(s=s1:Event[apply for a passport], p=p1:RequiredDocuments, o=o1:Documents)
Step2: Does this set of documents include the original identity card?
Action2: Deduce(left=o1, right=the original identity card, op=entailment)
Table 2: The cases of reasoning with logical form
@@ 2.3.3 Logical Form for Retrieval
In RAG, retrieval is achieved by calculating the similarity (e.g. cosine similarity) between the embeddings of the question and document chunks, where the semantic representation capability of embedding models plays a key role. This mainly includes a sparse encoder (BM25) and a dense retriever (BERT architecture pre-training language models). Sparse and dense embedding approaches capture different relevance features and can benefit from each other by leveraging complementary relevance information.
@@
For example, for the query "What documents are required to apply for a disability certificate at West Lake, Hangzhou?", the Retrieval Function could be represented as: "Retrieval(s=s1:Event[applying for a disability certificate], p=p1:RequiredDocuments, o=o1:Documents, s.location=West Lake, Hangzhou)". This approach leverages the establishment of different indices (sparse or dense) to facilitate precise searches or fuzzy searches as needed.
Furthermore, when structured knowledge in the form of SPO (Subject-Predicate-Object) cannot be retrieved using logical functions, alternative approaches can be employed. These include semi-structured retrieval, which involves using logical functions to search through chunks of information, and unstructured retrieval. The latter encompasses methods such as Retrieval-Augmented Generation (RAG), where sub-problems expressed in natural language are used to retrieve relevant chunks of text. This highlights the adaptability of the system to leverage different retrieval strategies based on the availability and nature of the information.
@@ 2.5.1 Natural Language Understanding
We fine-tuned six foundational models: qwen2, llama2, baichuan2, llama3, mistral, phi3, and used six understanding benchmarks recorded on OpenCompass for performance validation. The table below shows that the KAG-Model has a significant improvement in NLU tasks
@@ 2.5.2 Natural Language Inference
The NLI task is used to infer the semantic relations between given phrases. Typical NLI tasks include entity linking, entity disambiguation, taxonomy expansion, hypernym discovery, and text entailment. In the context of knowledge base question answering, due to the diversity and ambiguity of natural language expressions, as well as the subtle and different types of semantic connections between phrases, it often requires further alignment or retrieval of related information through NLI tasks based on NLU. As described in section 2.4, we categorize the key semantic relation in knowledge base applications into six types. Among these, relations such as isA, isPartOf and contains exhibit directional and distance-based partial order relations. During the reasoning process, it is crucial to accurately determine these semantic relations to advance towards the target answer. In traditional approaches, separate training of representation pre-training models and KG completion(KGC) models is often employed to reason about semantic relations. However, these KGC models tend to focus on learning graph structures and do not fully utilize the essential textual semantic information for semantic graph reasoning. LLMs possessing richer intrinsic knowledge, can leverage both semantic and structural information to achieve more precise reasoning outcomes. To this end, we have collected a high-quality conceptual knowledge base and ontologies from various domains, creating a conceptual knowledge set that includes 8,000 concepts and their semantic relations. Based on this knowledge set, we constructed a training dataset that includes six different types of conceptual reasoning instructions to enhance the semantic reasoning capabilities of a given base model, thereby providing semantic reasoning support for KAG.
Semantic reasoning is one of the core ability required in KAG process, we evaluate this ability of our model with two tasks, which is: downstream NLI tasks and general reasoning QA task, the result are as shown in Table 5, Table 6.
The evaluation results indicates that our KAG-Model demonstrates a significant improvement in tasks related with semantic reasoning: First, Table 6 shows that on the Hypernym Discovery task(which is consistent in form with the reasoning required in semantic enhanced indexing and retrieval.), our fine-tuned KAG-llama model outperforms Llama3 and ChatGPT-3.5 significantly. In addition, the better performance of our model on CMNLI, OCNLI and SIQA compared with Llama3 in Table 5 shows that our model has good capabilities in general logical reasoning.

### v2

@@ 2.2.3 Domain Knowledge Injection And Constraints
When openIE is applied to professional domains, irrelevant noise will be introduced. Previous researches[3, 5, 17] have shown that noisy and irrelevant corpora can significantly undermine the performance of LLMs. It is a challenge to align the granularity of extracted information and domain knowledge. The domain knowledge alignment capabilities in KAG include: 1) Domain term and concept injection. We use an iterative extraction approach, First, we store domain concepts and terms with description in KG storage. Second, we extract all instances in the document through openIE, then we perform vector retrieval to obtain all possible concept and term sets $E_{d}$. Finally, we add $E_{d}$ to the extraction prompt and perform another extraction to obtain a set $E_{d}^{a}$ that is mostly aligned with the domain knowledge. 2) Schema-constraint Extraction. In the vertical professional domains, the data structure between multiple documents in each data source such as drug instructions, physical examination reports, government affairs, online order data, structured data tables, etc. has strong consistency, and is more suitable for information extraction with schema-constraint, structured Extraction also makes it easier to do knowledge management and quality improvement. For detailed information about knowledge construction based on Schema-constraint, please refer to the SPG [^1] Official site of SPG: https://spg.openkg.cn/en-US and OneKE[18]. This article will not introduce it in detail. It is worth noting that, as shown in figure 2, for the same entity type, such as Person, we can pre-define properties and relations such as name, gender, placeOfBirth, (Person, hasFather, Person), (Person, hasFriend, Person), and can also extract tripples directly such as (Jay Chou, spgClass, Person), (Jay Chou, constellation, Capricorn), (Jay Chou, record company, Universal Music Group) through openIE. 3) Pre-defined Knowledge Structures By Document Type. Professional documents such as drug instructions, government affairs documents, and legal definitions generally have a relatively standardized document structure. Each type of document can be defined as an entity type, and different paragraphs are different properties of the entity. Taking government affairs as an example, we can pre-define the GovernmentAffair EntityType and properites such as administrative divisions, service procedures, required materials, service locations, and target groups. The divided chunks are the values of different properties. If the user asks "What materials are needed to apply for housing provident fund in Xihu District?", you can directly take out the chunk corresponding to property required materials to answer the question, avoiding the possible hallucinations caused by LLM re-generation.
@@ 2.2.4 Mutual indexing between text chunk vectors and knowledge structures
KAG’s mutual-indexing is a knowledge management and storage mechanism that conforms to the LLMFriSPG semantic representation. As is described in section 2.1, it includes four core data structures: 1) Shared Schemas are coarse-grained-types pre-defined as SPG Classes at project level, it includes EntityTypes, ConceptTypes, and EventTypes, they serve as a high-level categorization such as Person, Organization, GEOLocation, Date, Creature, Work, Event. 2) Instance Graph include all event and entity instances in $KG_{cs}$ and $KG_{fr}$. that is, instances constructed through openIE with schema-free or structured extraction with schema-constraint are both stored as instances in KG storage. 3) Text Chunks are special entity node that conforms to the definition of the Chunk EntityType. 4) Concept Graph is the core component for knowledge alignment. it consists of a series of concepts and concept relations, concept nodes are also fine-grained-types of instances. Through relation prediction, instance nodes can be linked to concept nodes to obtain their fine-grained semantic types. , and two storage structures: 1) KG Store. Store KG data structures in LPG databases, such as TuGraph, Neo4J. 2) Vector Store. Store text and vectors in a vector storage engine, such as ElasticSearch, Milvus, or the vector storage embedded in the LPG engine.
@@ 2.3 Logical Form Solver
In the process of solving complex problems, three key steps are involved: planning, reasoning and retrieval. Disassembling question is a planning process to determine the next problem to be tackled. Reasoning includes retrieving information based on the disassembled question, inferring the answer to the question according to the retrieved results, or re-disassembling the sub-question when the retrieved content cannot answer the question. Retrieval is to find the content that can be used as reference for the original question or the disassembled sub-question.
@@
Since interactions between different modules in traditional RAG are based on vector representations of natural language, inaccuracies often arise. Inspired by the logical forms commonly used in KGQA, we designed an executable language with reasoning and retrieval capabilities. This language breaks down a question into multiple logical expressions, each of which may include functions for retrieval or logical operations. The mutual indexing described in Section 2.2 makes this process possible. Meanwhile, we designed a multi-turn solving mechanism based on reflection and global memory, inspired by ReSP[19]. The KAG solving process, as referenced in Figure 6 and Algorithm 17, first decomposes the current question $query_{cur}$ into a list of subquestions $lf_{list}$ represented in logical form, and performs hybrid reasoning to solve them. If an exact answer can be obtained through multi-hop reasoning over structured knowledge, it returns the answer directly. Otherwise, it reflects on the solution results: storing the answers and retrieval results corresponding to $lf_{list}$ in global memory and determining whether the question is resolved. If not, it generates supplementary questions and proceeds to the next iteration. Section 2.3.1, 2.3.2 and 2.3.3 introduce logical form function for planning, logical form for reasoning and logical form for retrieval respectively. In general, the proposed logical form language has the following three advantages:
  • The use of symbolic language enhances the rigor and interpretability of problem decomposition and reasoning.
  • Make full use of LLMFriSPG hierarchical representation to retrieve facts and texts knowledge guided by the symbolic graph structure
@@
Table 13 illustrates a multi-round scenario consistent with pseudocode 17. Although first round the exact number of plague occurrences couldn’t be determined, but we can extracted information indicates: "Venice, the birthplace of Antonio Vivaldi, experienced the devastating Black Death, also known as the Great Plague. This pandemic caused by Yersinia pestis led to 75 to 200 million deaths in Eurasia, peaking in Europe from 1347 to 1351. The plague brought significant upheavals in Europe. Although specific occurrence records in Venice aren’t detailed, it’s clear the city was impacted during the mid-14th century.". As is shown in Table 13,After two iterations, the answer determined is: 22 times.
2.3.1 Logical Form Planning
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Retrieval}(s=s_{i}:type[name],p=p_{i}:edge,o=o_{i}:type[name],$
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Sort}(A,direction=min|max,limit=n)$
@@
$math_{i}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Math}(expr)$,
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Deduce}(left=A,right=B,op=entailment|greater|less|equal)$
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Output}(A,B,...)$
@@
Retrieval. According to the the knowledge or information retrieved from SPO, s, p, o should not repeatedly appear multiple times in the same expression. Constraints can be applied to the s, p, o for querying. For multi-hop queries, multiple retrievals are required. When the current variable refers to a previously mentioned variable, the variable name must be consistent with the referenced variable name, and only the variable name needs to be provided. The knowledge type and name are only specified during the first reference.
Sort. Sort the retrieved results. $A$ is the variable name for the retrieved subject-predicate-object(SPO) ($s_{i}$, $o_{i}$, or $s.prop$, $p.prop$, $o.prop$). $direction$ specifies the sorting direction, where $direction=min$ means sorting in ascending order and $direction=max$ means sorting in descending order. $limit=n$ indicates outputting the topN results.
Math. Perform mathematical calculations. $expr$ is in LaTeX syntax and can be used to perform calculations on the retrieved results (sets) or constants. $math_{i}$ represents the result of the calculation and can be used as a variable name for reference in subsequent actions.
@@ 2.3.2 Logical Form for Reasoning
When the query statement represented by natural language is applied to the search, the logic is often fuzzy, such as "find a picture containing vegetables or fruits" and "find a picture containing vegetables and fruits". Whether text search or vector search is used, the similarity between the two queries is very high, but the corresponding answers are quite different. The same is true for problems involving logical reasoning processes such as and or not, and intersection differences. To this end, we use logical form to express the question, so that it can express explicit semantic relations. Similar to IRCOT, we decompose complex original problem and plan out various execution actions such as multi-step retrieval, numerical reasoning, logical reasoning, and semantic deduce. Each sub-problem is expressed using logical form functions, and dependencies between sub-questions are established through variable references. The inference resolution process for each sub-question is illustrated as Algorithm 11. In this process, the GraphRetrieval module performs KG structure retrieval according to the logical form clause to obtain structured graph results. Another key module, HybridRetrieval, combining natural language expressed sub-problems and logical functions for comprehensive retrieval of documents and sub-graph information. To understand how logical functions can be utilized to reason about complex problems, refer to the following examples as Table 14.
@@ 2.3.3 Logical Form for Retrieval
In naive RAG, retrieval is achieved by calculating the similarity (e.g. cosine similarity) between the embeddings of the question and document chunks, where the semantic representation capability of embedding models plays a key role. This mainly includes a sparse encoder (BM25) and a dense retriever (BERT architecture pre-training language models). Sparse and dense embedding approaches capture different relevance features and can benefit from each other by leveraging complementary relevance information.
@@
For example, for the query "What documents are required to apply for a disability certificate at West Lake, Hangzhou?", the retrieval function could be represented as: "Retrieval(s=s1:Event[applying for a disability certificate], p=p1:support_chunks, o=o1:Chunk, s.location=West Lake, Hangzhou)". This approach leverages the establishment of different indices (sparse or dense) to facilitate precise searches or fuzzy searches as needed.
Furthermore, when structured knowledge in the form of SPO cannot be retrieved using logical functions, alternative approaches can be employed. These include semi-structured retrieval, which involves using logical functions to search through chunks of information, and unstructured retrieval. The latter encompasses methods such as Retrieval-Augmented Generation (RAG), where sub-problems expressed in natural language are used to retrieve relevant chunks of text. This highlights the adaptability of the system to leverage different retrieval strategies based on the availability and nature of the information.
@@ 2.5.1 Natural Language Understanding
We fine-tuned six foundational models: qwen2, llama2, baichuan2, llama3, mistral, phi3, and used six understanding benchmarks recorded on OpenCompass for performance validation. The table 3 shows that the KAG-Model has a significant improvement in NLU tasks.
@@ 2.5.2 Natural Language Inference
The NLI task is used to infer the semantic relations between given phrases. Typical NLI tasks include entity linking, entity disambiguation, taxonomy expansion, hypernym discovery, and text entailment. In the context of knowledge base Q&A, due to the diversity and ambiguity of natural language expressions, as well as the subtle and different types of semantic connections between phrases, it often requires further alignment or retrieval of related information through NLI tasks based on NLU. As described in section 2.4, we categorize the key semantic relation in knowledge base applications into six types. Among these, relations such as isA, isPartOf and contains exhibit directional and distance-based partial order relations. During the reasoning process, it is crucial to accurately determine these semantic relations to advance towards the target answer. In traditional approaches, separate training of representation pre-training models and KG completion(KGC) models is often employed to reason about semantic relations. However, these KGC models tend to focus on learning graph structures and do not fully utilize the essential textual semantic information for semantic graph reasoning. LLMs possess richer intrinsic knowledge, and can leverage both semantic and structural information to achieve more precise reasoning outcomes. To this end, we have collected a high-quality conceptual knowledge base and ontologies from various domains, creating a conceptual knowledge set that includes 8,000 concepts and their semantic relations. Based on this knowledge set, we constructed a training dataset that includes six different types of conceptual reasoning instructions to enhance the semantic reasoning capabilities of a given base model, thereby providing semantic reasoning support for KAG.
Semantic reasoning is one of the core ability required in KAG process, we use NLI tasks and general reasoning Q&A tasks to evaluate the ability of our model, the results are as shown in Table 4 and Table 5. The evaluation results indicates that our KAG-Model demonstrates a significant improvement in tasks related with semantic reasoning: First, Table 5 shows that on the Hypernym Discovery task(which is consistent in form with the reasoning required in semantic enhanced indexing and retrieval.), our fine-tuned KAG-llama model outperforms Llama3 and ChatGPT-3.5 significantly. In addition, the better performance of our model on CMNLI, OCNLI and SIQA compared with Llama3 in Table 4 shows that our model has good capabilities in general logical reasoning.